On dissolving 12 g urea in 2L of 0.05 M aqueous solution of urea , which of the following concentrations is/are true for the resulting solution : (a) molarity =0.15 M (b) %w/v=9% (c) strength =15g/L (d) X urea=0.3

Grade:10

1 Answers

Khimraj
5 years ago
initial mole of urea = 0.05*2 = 0.1
final mole = 0.1 + 12/60 = 0.3
weight of urea = 0.3*60 = 18 g
​(a) Molarity = 0.3/2 = 0.15M
(b) %w/V = (18/2000)*100 = 0.9%
(c) stregth = 18/2 = 9g/L
(d) X = 0.3/0.3 = 1
Hope it clears...................